Write the equation of the line parallel to the given line and passing through the givern point 6 through (3,4)
Answer #1

Equation y=6 represents the slope of line m=0

Since the parallel line passing through (3,4), hence the point intercept form of line satisfy the given point.

Equation is as : (y-y1) = m (x-x1)

(y-4) = 0(x-3)

Y=4 ( line parallel to the given line and passion through given point)